OpenStudy (anonymous):

Two garden plots are to have the same area. One is square and one is rectangular. The rectangular plot is 2 meters wide and 8 meters long. How large is one side of the square garden plot in meters?

OpenStudy (anonymous):

yeah so here we first need to find the area of the rectangluar plot can u find it

OpenStudy (anonymous):

u there ??

OpenStudy (anonymous):

oh yea im here

OpenStudy (anonymous):

is it 16 meters?

OpenStudy (anonymous):

yeah it is 16 m ^2

OpenStudy (anonymous):

so noww u know the area of the square is sides^2

OpenStudy (anonymous):

let the side of the square plot be x then x^2 = 16m^2

OpenStudy (anonymous):

>.<

OpenStudy (anonymous):

x^2 = 16m^2

OpenStudy (anonymous):

so x should be \[\sqrt{16}\] which is 4

OpenStudy (anonymous):

so the side is 4

OpenStudy (anonymous):

it is thanks @Smita12

OpenStudy (anonymous):

:DDD happy to help u